OpenStudy (anonymous):

quick question is 1 - tan^2 x/1 + tan^2x = cos2x? Why is it so if yes?

OpenStudy (anonymous):

@rational

OpenStudy (rational):

write tan^2 as sin^2/cos^2 and keep simplifying

OpenStudy (anonymous):

oh okay sorry.. another one is cos^2x - sin^2x = 1 - tan x?

OpenStudy (rational):

\[\large \begin{align} \dfrac{1-\tan^2x}{1+\tan^2x}&=\dfrac{1-\frac{\sin^2x}{\cos^2x}}{1+\frac{\sin^2x}{\cos^2x}}\\~\\ &= \dfrac{\cos^2x-\sin^2x}{\cos^2x+\sin^2x}\\~\\ &=\cdots \end{align}\]

OpenStudy (anonymous):

I did not write down cos^2 x - sin^x/cos^x + sin^x formula :(

OpenStudy (anonymous):

Is there any identity for root1+sinx + root1-sinx?

OpenStudy (rational):

\[\large \sqrt{1+\sin x} + \sqrt{1-\sin x}\] like that?

OpenStudy (anonymous):

yeah!

OpenStudy (rational):

it wont simplify i guess http://www.wolframalpha.com/input/?i=alternate+forms++ \sqrt{1%2B\sin+x}+%2B+\sqrt{1-\sin+x}

OpenStudy (anonymous):

hang on lemme use the equation thing on openstudy.. sec

OpenStudy (rational):

kk